# SoK: A Classification for AI-driven Personalized Privacy Assistants

### v2

@@ A.1.3. Study Records
  • The presence of a user study
  • Extent of evaluation – scale of validation activity that is measured
@@
  • Quality assessment and critical appraisal of the studies that have validated or evaluated the AI-driven PPA
@@
  • Features of users control over decisions (initially guided by EU consent requirements)
Types of contributions
Inspired by Kuhrmann et al. and Shaw , we classified publications by their types of contributions (i.e., multiple choice) according to the following:
Model:
Representation of observed reality by concepts after conceptualization.
Theory:
Construct of cause-effect relationships.
Framework:
Frameworks/methods (related to automated privacy decisions).
Guidelines:
List of advice.
Lessons learned:
Set of outcomes from obtained results.
Advice:
Recommendation (from opinion).
Tool:
A tool to automate privacy decisions.
Data synthesis
We collated and summarized results into classification tables. We composed a narrative synthesis for papers meeting our inclusion criteria.
Preliminary components of the data synthesis:
  • Overall identification and classification of AI-driven PPAs in published research
  • Classification tables presenting features of AI-driven PPAs used to coherently organize the solutions surveyed
  • Comparison analysis – based on the features of AI-driven PPAs–, and narrative synthesis
A.2. Further Results from Data Charting and Critical Appraisal
@@
United States 13 China 2
@@
Table 4. Countries of affiliation of authors of selected papers.
@@
2018 5
Table 5. Number of publications per year.
@@
[b] Type of contribution Publication Accuracy User study Critical appraisal Framework Tool Model Theory Lessons learned Advice Xie et al. 68% Online user experiment $\alpha$bullet $\bullet$ Apolinarski et al. No $\bullet$ $\bullet$ Hirschprung et al. Online qualitative survey D-, very low (55%) $\bullet$ $\bullet$ Squicciarini et al. 92.53% Cross sectional study  [^15] A survey-based study and a direct user evaluation D, very low (55%) $\bullet$ $\bullet$ Liu et al. 78.7% Randomized controlled studies  [^16] Two surveys A, high (90%) $\bullet$ $\bullet$ Albertini et al. Cross-sectional study D, very low (55%) $\bullet$ Dong et al. 89,8% F1 Case studies $\alpha$bullet $\bullet$ Baarslag et al. Randomized controlled study  [^17] Between-participants design A, high (90%) $\bullet$ $\bullet$ $\bullet$ Fogues and Murukannaiah Around 50% Online survey $\alpha$bullet Zhong et al. 79% Survey $\alpha$bullet Misra and Such 91.8% Non-controlled before-after study  [^18] Online survey C, limited (70%) $\bullet$ Nakamura et al. 85% Cross-sectional study  [^19] Online questionnaire D, very low (55%) $\bullet$ Olejnik et al. More than 80% Yes, for data collection $\alpha$bullet $\bullet$ Das et al. No $\bullet$ Tan et al. 95%  [^20] For privacy leakage detection (not to be confused with preferences detection) No $\bullet$ Wijesekera et al. 95% Experience Sampling Method D, low (60%) $\bullet$ $\bullet$ Yu et al. Cross-sectional study  [^21] To measure the interpretability of the approaches D, very low (55%) $\bullet$ $\bullet$ Bahirat et al. 81.54% No $\bullet$ Klingensmith et al. No $\bullet$ Barbosa et al. 86.8%  [^22] AUC of binary allow/deny for a given scenario Survey $\alpha$bullet $\bullet$ Alom et al. [b] Up to 72.2% (satisfaction) Cross-sectional study  [^23] User satisfaction D, very low (55%) $\bullet$ Alom et al. [a] 96.4% and 94.5%  [^24] Accuracy based on appreciation of evaluators Yes, for labeling and evaluation $\alpha$bullet Kasaraneni and Thomas No $\bullet$ $\bullet$ Kaur et al. [a] No $\bullet$ Botti-Cebriá et al. –  [^25] The accuracy presented is for the right category of data No $\bullet$ Kokciyan and Yolum No $\bullet$ Sanchez et al. 84.74% Online survey to build their dataset $\alpha$bullet Kaur et al. [b] No $\bullet$ Lobner et al. 83.33%  [^26] With interpretability of the results Survey $\alpha$bullet Filipczuk et al. 65%  [^27] On average, but seems higher in specific case Between-subject experimental design C, limited (70%) $\bullet$ $\bullet$ Hirschprung and Alkoby Cross-sectional study D, low (60%) $\bullet$ $\bullet$ Kökciyan and Yolum Between 41 and 92%  [^28] Depends on several parameters No $\bullet$ Ulusoy and Yolum Around 75%  [^29] Difficult to assess because they measure utility of decisions in a simulated setting No $\bullet$ Zhan et al. 74% No $\bullet$ Brandão et al. Between 82 and 88% Field study (cross-sectional study) D, low (60%) $\bullet$ Mendes et al. 92% Cross-sectional study  [^30] Field study D, low (60%) $\bullet$ $\bullet$ $\bullet$ Shanmugarasa et al. 92.62% Cross-sectional study  [^31] Field study D, low (60%) $\bullet$ $\bullet$ Ayci et al. 89% No $\bullet$ $\bullet$ Serramia et al. 3.78/5  [^32] Acceptability rate, not accuracy Cross-sectional study  [^33] To measure the level of comfort of the norms inferred D, very low (55%) $\bullet$ $\bullet$

### v4

@@ III-B4 Data Extraction and Analysis
- The presence of a user study, and performed its critical appraisal (see below).
- Extent of evaluation, a scale of validation activity that is measured.
@@
- Quality assessment and critical appraisal of the studies that have validated or evaluated the AI-driven PPA.
@@
- Features of user’s control over decisions (initially guided by EU consent requirements).
It is also worth noting that we classified publications by their types of contributions according to the following items proposed by Kuhrmann et al. [20] and Shaw [21] (i.e., all that apply): (i) model, as a representation of observed reality by concepts after conceptualization; (ii) theory, as a construct of cause-effect relationships; (iii) framework, including frameworks/methods (related to automated privacy decisions); (iv) guidelines, as a list of advice; (v) lessons learned, as a set of outcomes from obtained results; (vi) advice, as recommendations (from opinion); and, (vii) tool, as tools to automate privacy decisions.
TABLE III: This table informs the type of contribution, informing whether the surveyed solution presents a framework, a tool (i.e. with an implementation), a model, lessons learned, or advice.
Type of contribution
Year Publication
Framework
Tool
Model
Theory
Lessons learned
Advice
2014 Xie et al. [22] $\bullet$ $\bullet$
2015 Apolinarski et al. [23] $\bullet$ $\bullet$
2015 Hirschprung et al. [24] $\bullet$ $\bullet$
2015 Squicciarini et al. [25] $\bullet$ $\bullet$
2016 Liu et al. [26] $\bullet$ $\bullet$
2016 Albertini et al. [27] $\bullet$
2016 Dong et al. [28] $\bullet$ $\bullet$
2017 Baarslag et al. [29] $\bullet$ $\bullet$ $\bullet$
2017 Fogues et al. [30] $\bullet$
2017 Zhong et al. [31] $\bullet$
2017 Misra et al. [32] $\bullet$
2017 Nakamura et al. [33] $\bullet$
2017 Olejnik et al. [34] $\bullet$ $\bullet$
2018 Das et al.[35] $\bullet$
2018 Tan et al. [36] $\bullet$
2018 Wijesekera et al. [37] $\bullet$ $\bullet$
2018 Yu et al. [38] $\bullet$ $\bullet$
2018 Bahirat et al. [39] $\bullet$
2018 Raber et al. [40] $\bullet$ $\bullet$ $\bullet$
2019 Klingensmith et al. [41] $\bullet$
2019 Barbosa et al. [42] $\bullet$ $\bullet$
2019 Alom et al. [43] $\bullet$
2019 Alom et al. [44] $\bullet$
2020 Kasaraneni et al. [45] $\bullet$ $\bullet$
2020 Kaur et al. [46] $\bullet$
2020 Botti-Cebria et al. [47] $\bullet$
2020 Kökciyan et al. [48] $\bullet$
2020 Sanchez et al. [49] $\bullet$
2021 Kaur et al. [50] $\bullet$
2021 Lobner et al. [51] $\bullet$
2022 Filipczuk et al. [52] $\bullet$ $\bullet$
2022 Hirschprung et al. [53] $\bullet$ $\bullet$
2022 Kökciyan et al. [54] $\bullet$
2022 Ulusoy et al. [55] $\bullet$
2022 Zhan et al. [56] $\bullet$
2022 Brandão et al. [57] $\bullet$
2022 Mendes et al. [58] $\bullet$ $\bullet$ $\bullet$
2022 Shanmugarasa et al. [59] $\bullet$ $\bullet$
2023 Ayci et al. [60] $\bullet$ $\bullet$
2023 Serramia et al. [61] $\bullet$ $\bullet$
2024 Wang et al. [62] $\bullet$
Although we attempted to extract as much data from the studies as possible using a DEF, we found that, during the data analysis process, there was a need to further categorize studies across other facets. For instance, additional information was compiled in the spreadsheets, such as a high-level categorization of certain fields (i.e., the type of AI used) or a critical appraisal of the user studies presented in the selected papers, although only when those user studies were used to evaluate the AI-driven PPA, and not when they were used for data collection to build datasets.
This collection of facets created during the study design and data analysis processes forms the basis of the work’s final classification scheme, presented as part of the main results. All authors were involved in the data analysis process and the definition of facets that further classify studies on the topic.
For the critical appraisal, we used the CAT (Critically Appraised Topic) Manager App of CEBMA (the Center for Evidence-Based Management) [63], which provides a practical yet rigorous approach to evaluate studies based on objective criteria. It helps determine a study’s trustworthiness regarding cause and effect questions. Once a study is evaluated, the possible outcomes are: Very high (A+), High (A), Moderate (B), Limited (C), Low (D), or Very low (D-).
By definition, an AI-driven PPA leverages AI techniques. We therefore collected information about the Type of AI used. AI models rely on data for training and decision-making. As such, we extracted the source of data. During the adjustment of the DEF, we observed that AI-driven PPAs are usually designed for a specific system context, and for one or several types of decision. Connected to the system context, we extracted the choice of architecture of the implementation (if any) to analyze the trust implications.
We also collected the methods for an empirical assessment, presence, and quality of user studies, or the means used to measure the accuracy, to gain insights on eventual benchmarks of AI-driven PPAs. Studies can be classified as evaluation or validation research, as proposed by Wieringa et al. [64]. An evaluation works in real-world practice and is implementing/deploying the solution or testing in an actual project with real test users, such as real case studies and realistic user testing of prototypes/systems. A validation is a limited illustrative or hypothetical “case study” or “use case” performed as a lab experiment. In general practice, prototypes are often validated by cross-sectional studies.
Finally, initially guided by legal requirements for consent and the exercise of data subject rights under the GDPR (although eventually, no paper considers consent), we extracted what became user control over decisions.
III-C Reporting the Review
Based on the data analysis, a whole coherent narrative was written by the research team, i.e., this SLR article, conveying all the results, our interpretation of the main findings, and identifying research gaps. This synthesis on AI-driven PPAs is thus reported in the following sections.
IV Summary of Data Charting Results
This section provides an overview of the quantitative insights generated through the data charting process (e.g., publications per year, citations, types of decisions). The main findings related to the critical appraisals are also introduced in this section. It is worth noting, nonetheless, that the classification features are further detailed in the following Section V.
Among the 41 papers surveyed, we tallied 15 different countries for the authors’ affiliations (see Table IV), with the USA and UK leading in numbers. About 55% ($n=22$) of the selected publications were published from 2019 to 2024, with the year 2021 being the most productive with 8 publications (see Table V). At the time of data collection, papers were cited between 0 and 275 times with an average of 38.56 citations, a median of 14, and a standard deviation of 59, indicating a power law distribution of the citation count. The most cited papers are Liu et al. [26] ($n=275$), Yu et al. [38] ($n=199$), and Squicciarini et al. [25] ($n=130$) (numbers at the time of data collection).
TABLE IV: Countries of affiliation of authors of selected papers.
@@ IV Summary of Data Charting Results
United States 14 China 2
@@
TABLE V: Number of publications per year.
@@
2018 5 2024 1
As shown in Table VI, regarding the sources of data used by the AI-driven PPAs, context data, attitudinal data, and metadata were the most prevalent. We observed a relatively balanced distribution when it comes to the types of decisions (between 12 and 15 for each type) and the system contexts (between 11 and 13, with two outliers for Cloud and Intelligent retail store). For the types of AI systems that we were able to classify, most models were deemed non-intrinsically transparent (NIT, $n=14$), followed by transparent (T, $n=8$) and partially transparent (PT, $n=4$) models. Note also that Das et al. [35] did not specify the type of AI used in their paper, we were therefore unable to categorize their solution in that respect (under Type of AI used).
In Table III, the publications were also classified by their types of contributions, according to the categories proposed by Kuhrmann et al. [20] and Shaw [21]. We observed a prevalence of models ($n=24$) and tools ($n=21$), followed by frameworks ($n=9$). Nonetheless, these models, tools, and frameworks lack empirical assessment, an issue further analyzed in Section V-F.
Finally, the results of our critical appraisal can be found in Table VII. Out of the 41 publications, only 15 presented a user study, i.e., qualitative research that is suitable to be critically appraised. In terms of quality, they mostly scored “low” or “very low” ($n=9$) according to the CEBMA checklist. Exceptionally, only the studies of Liu et al. [26] and Baarslag et al. [29] were appraised as of high quality. This suggests that the empirical evidence around AI-driven PPAs remains incipient, and existing solutions can be further tested in real-world settings, a challenge (or opportunity) that is discussed in Section VI.
V Classification for AI-driven PPAs
We provide in this section a classification for AI-driven PPAs as the main contribution of this SLR. Summarized in Figure 2, the classification comprises several dimensions, i.e., features typically considered in the design of such an assistant (see also Tables VI and VII). These dimensions are the type of decision (Section V-A), the type of AI (Section V-B) and the source of data (Section V-C) used in the decision, the system context (Section V-D), the choice architecture of its eventual implementation (Section V-E), the empirical assessment (Section V-F), and the extent to which users have control over the decisions (Section V-G).
TABLE VI: Summary table of our classification, part 1.   It presents the mandatory features of AI-driven PPAs, namely the type of decision, the type of AI used (for which we specified whether the classification model is Transparent (T), Not-Inherently Transparent (NIT), or Partially Transparent (PT) because several models are used), the type of source of data, and the system context (note that IRS stands for Intelligent Retail Store).   An empty field signifies that the solution does not exhibit the characteristic (e.g., does not consider Y type of decision).
Type of decision Type of AI used Type of source of data System context
Year Publication
Permissions
Preferences
Data sharing
Classification
Clustering
Rule-based
Logic-based
Reinforcement
LLM
Context
Attitudinal data
Metadata
Data type
Content of data
Behavioral data
Mobile apps
Social media
IoT
Cloud
IRS
2014 Xie et al. [22] $\bullet$ NIT $\bullet$
2015 Apolinarski et al. [23] $\bullet$ NIT $\bullet$ $\bullet$
2015 Hirschprung et al. [24] $\bullet$ $\bullet$ $\bullet$ $\bullet$
2015 Squicciarini et al. [25] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2016 Liu et al. [26] $\bullet$ T $\bullet$ $\bullet$ $\bullet$
2016 Albertini et al. [27] $\bullet$ $\bullet$ $\bullet$ $\bullet$
2016 Dong et al. [28] $\bullet$ T $\bullet$ $\bullet$ $\bullet$
2017 Baarslag et al. [29] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2017 Fogues et al. [30] $\bullet$ PT $\bullet$ $\bullet$ $\bullet$
2017 Zhong et al. [31] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2017 Misra et al. [32] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2017 Nakamura et al. [33] $\bullet$ NIT $\bullet$ $\bullet$
2017 Olejnik et al. [34] $\bullet$ T $\bullet$ $\bullet$ $\bullet$
2018 Das et al.[35] $\bullet$ $\bullet$ $\bullet$
2018 Tan et al. [36] $\bullet$ T $\bullet$ $\bullet$
2018 Wijesekera et al. [37] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$ $\bullet$
2018 Yu et al. [38] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2018 Bahirat et al. [39] $\bullet$ T $\bullet$ $\bullet$
2018 Raber et al. [40] $\bullet$ T $\bullet$ $\bullet$ $\bullet$
2019 Klingensmith et al. [41] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2019 Barbosa et al. [42] $\bullet$ PT $\bullet$ $\bullet$ $\bullet$
2019 Alom et al. [43] $\bullet$ PT $\bullet$ $\bullet$
2019 Alom et al. [44] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2020 Kasaraneni et al. [45] $\bullet$ T $\bullet$ $\bullet$ $\bullet$
2020 Kaur et al. [46] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$ $\bullet$
2020 Botti-Cebria et al. [47] $\bullet$ PT $\bullet$ $\bullet$
2020 Kökciyan et al. [48] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2020 Sanchez et al. [49] $\bullet$ $\bullet$ $\bullet$ $\bullet$
2021 Kaur et al. [50] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2021 Lobner et al. [51] $\bullet$ T $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Filipczuk et al. [52] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Hirschprung et al. [53] $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Kökciyan et al. [54] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Ulusoy et al. [55] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Zhan et al. [56] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2022 Brandão et al. [57] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2022 Mendes et al. [58] $\bullet$ NIT $\bullet$ $\bullet$
2022 Shanmugarasa et al. [59] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2023 Ayci et al. [60] $\bullet$ NIT $\bullet$ $\bullet$ $\bullet$
2023 Serramia et al. [61] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
2024 Wang et al. [62] $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
TABLE VII: Summary table of our classification, part 2.   It presents the optional features of AI-driven PPAs, such the Architecture, under which we denote with “–” when the criterion is not applicable (no implementation/tool is presented) and when the solution presents an implementation, but the paper did not specify enough information to infer its architecture.   For user control over decisions, we specify the elements present to inform users under Informed (type of Data, Purpose, Controller).   It also presents the accuracy (if any) of the predictions (see Section V-F); the type of user control over decisions; the presence or not of a user study, and the type of user study if applicable; and the results of our critical appraisal (see Section III-B4).
@@ V Classification for AI-driven PPAs
[b] Architecture User control over decision Year Publication Local Remote Federated Informed Semi-automated Specific Revoke Accuracy User study Critical appraisal 2014 Xie et al. [22] No Yes Yes a No 68% Online user experiment $\alpha$ 2015 Apolinarski et al. [23] $\bullet$ D Yes Yes No No 2015 Hirschprung et al. [24] D No b Yes No Online qualitative survey D-, very low (55%) 2015 Squicciarini et al. [25] D Yes Yes No 92.53% Cross sectional study c D-, very low (55%) 2016 Liu et al. [26] ? ? D, P Yes Yes Yes 78.7% Randomized controlled studies d A, high (90%) 2016 Albertini et al. [27] $\bullet$ D Yes No No Cross-sectional study D, very low (55%) 2016 Dong et al. [28] 89,8% F1 Case studies $\alpha$ 2017 Baarslag et al. [29] $\bullet$ $\bullet$ Unclear Yes Yes No Randomized controlled study e A, high (90%) 2017 Fogues et al. [30] $\bullet$ No No No No Around 50% Online survey $\alpha$ 2017 Zhong et al. [31] 79% Survey $\alpha$ 2017 Misra et al. [32] $\bullet$ D Yes Yes No 91.8% Non-controlled before-after study f C, limited (70%) 2017 Nakamura et al. [33] 85% Cross-sectional study g D, very low (55%) 2017 Olejnik et al. [34] $\bullet$ No Yes Yes No More than 80% Yes, for data collection $\alpha$ 2018 Das et al.[35] $\bullet$ Yes It depends Yes No No 2018 Tan et al. [36] $\bullet$ No No h Yes No 95% i No 2018 Wijesekera et al. [37] $\bullet$ $\bullet$ D, C Yes Yes Yes 95% Interrupted time series study (ESM) B, moderate (80%) 2018 Yu et al. [38] Cross-sectional study j D, very low (55%) 2018 Bahirat et al. [39] D, P k It depends It depends No 81.54% No 2018 Raber et al. [40] D Yesl Yes Yes 70% Non-controlled before-after study m C, limited (70%) 2019 Klingensmith et al. [41] $\bullet$ $\bullet$ D Not always Yes No No 2019 Barbosa et al. [42] 86.8% n Survey $\alpha$ 2019 Alom et al. [43] Up to 72.2% (satisfaction) Cross-sectional study o D, very low (55%) 2019 Alom et al. [44] 96.4% and 94.5% p Yes, for labeling and evaluation $\alpha$ 2020 Kasaraneni et al. [45] $\bullet$ D Yes Yes No No 2020 Kaur et al. [46] No 2020 Botti-Cebria et al. [47] $\bullet$ D Yes Yes No – q No 2020 Kökciyan et al. [48] No 2020 Sanchez et al. [49] Unclear Yes Yes No 84.74% Online survey to build their dataset $\alpha$ 2021 Kaur et al. [50] No 2021 Lobner et al. [51] 83.33% r Survey $\alpha$ 2022 Filipczuk et al. [52] $\bullet$ $\bullet$ D Yes Yes No 65% s Non-controlled before-after study t C, limited (70%) 2022 Hirschprung et al. [53] Cross-sectional study D, low (60%) 2022 Kökciyan et al. [54] No It depends Yes No Between 41 and 92% u No 2022 Ulusoy et al. [55] Around 75% v No 2022 Zhan et al. [56] 74% No 2022 Brandão et al. [57] $\bullet$ Between 82 and 88% Field study $\alpha$ 2022 Mendes et al. [58] $\bullet$ No No No No 92% Field study $\alpha$ 2022 Shanmugarasa et al. [59] $\bullet$ No Yes Yes No 92.62% Cross-sectional study w D, very low (55%) 2023 Ayci et al. [60] $\bullet$ No Yes Yes No 89% No 2023 Serramia et al. [61] $\bullet$ No Yes No No 3.78/5 x Cross-sectional study y D, very low (55%) 2024 Wang et al. [62] $\bullet$ No No Yes No No
  • a
    Only location
  • b
    Not necessarily, depends on what they call the Configuration Options
  • c
    A survey-based study and a direct user evaluation
  • d
    Two surveys
  • e
    Between-participants design
  • f
    Online survey
  • g
    Online questionnaire
  • h
    Not by default, they have a sort of ‘user settings’ for expert users
  • i
  • j
    To measure the interpretability of the approaches
  • k
    Not consistently
  • l
    Based on the current data collection
  • m
    Online survey
  • n
    AUC of binary allow/deny for a given scenario
  • o
    User satisfaction
  • p
    Accuracy based on appreciation of evaluators
  • q
    The accuracy presented is for the right category of data
  • r
    With interpretability of the results
  • s
    On average, but seems higher in specific case
  • t
    Between-subject experimental design
  • u
    Depends on several parameters
  • v
  • w
    Online survey
  • x
    Acceptability rate, not accuracy
  • y
    To measure the level of comfort of the norms inferred